One number is 2 times a first number. A third number is 100 more than the first number. If the sum of the three numbers is 456, find the numbers.
Answer by TimothyLamb(4379)

y = 2x
z = x + 100
x + y + z = 456
---
put the system of linear equations into standard form
---
2x - y = 0
x - z = -100
x + y + z = 456
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 89
y = 178
z = 189
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---